# DeToNATION: Decoupled Torch Network-Aware Training on Interlinked Online Nodes

### v3

@@ Abstract
Training large neural network models requires extensive computational resources, often distributed across several nodes and accelerators. Recent findings suggest that it may be sufficient to only exchange the fast moving components of the gradients, while accumulating momentum locally (Decoupled Momentum, or DeMo). However, DeMo assumes that models fit on a single accelerator. We relax this assumption and introduce FlexDeMo, whereby nodes fully shard model parameters locally between different accelerators, while inter-node communication is reduced by synchronizing only fast-moving components instead of the full gradients – resulting in a hybrid sharded data parallel training strategy. We further introduce a framework, denoted as DeToNATION, that generalizes DeMo, FlexDeMo, and other popular distributed training schemes such as DiLoCo – introducing new variations of replication schemes and challenging choices made in DeMo. Our results across language and vision domains show that FlexDeMo attains similar validation loss as hybrid sharded data parallel training employing AdamW and full gradient synchronization, while being substantially faster. FlexDeMo is thus a promising distributed training scheme for the largest machine learning models.
Codehttps://github.com/schneiderkamplab/DeToNATION/
@@ Introduction
Training large deep neural networks (DNNs) induces large amounts of network traffic in the form of gradients that are transmitted between accelerators, typically requiring expensive localized high-throughput network setups on high-performance computing clusters. The network throughput increasingly becomes a bottleneck, as the number of accelerator nodes participating in the training increases and the general network congestion increases. Recent work demonstrates that synchronizing the full optimizer state is not always necessary for state-of-the-art results through decoupling momentum updates and allowing controlled divergence within each rank by carefully controlled inter-accelerator communication termed Decoupled Momentum optimization (DeMo) (Peng, Quesnelle, and Kingma 2024).
DeMo presents a viable strategy for distributed training with reduced gradient communication and, thus, enables distributed data parallel (DDP) training of DNNs with relatively low network bandwidth requirements. However, this strategy comes with several caveats. First, relying on DDP implies the constraint that the DNN model and optimizer states must fit within the memory capacity of each accelerator. This severely limits the applicability for training large models such as state-of-the-art large language models, which commonly do not fit within the memory of a single accelerator. Second, DeMo inherently relies on a distributed gathering operation whose bandwidth requirements scale linearly with the number of accelerators. Third, DeMo leaves quite a lot of open questions regarding the replication-scheme and hyperparameters, such as chunk-size, TopK and employing the sign-function.
@@
Instead of synchronizing the full gradients between the nodes as in extant hybrid sharding strategies, we compress, gather, and decompress selected relevant fast-moving momentum components following the approach that the DeMo optimizer has introduced for DDP training (Peng, Quesnelle, and Kingma 2024). This relaxes constraints regarding the accelerator memory requirements while simultaneously reducing inter-node bandwidth requirements. We further investigate hyper-parameters of DeMo and challenge the fast-moving momenta-based replication scheme by introducing Striding, Random and the prior proposed DiLoCo schemes. This addresses the three caveats of DeMo mentioned above. First, FlexDeMo allows for decoupled momentum training of models that do not fit into the memory of a single accelerator but fit into the combined memory of the accelerators of one node. Second, FlexDeMo replicates between nodes rather than accelerators, effectively reducing the bandwidth requirements of the distributed gathering operation, which now scales linearly with the number of nodes rather than the number of accelerators. Third, we justify our choice of hyper-parameters by conducting experiments.
We validate our results on the T5, ViT and OLMo models in machine translation, image classification and causal language modeling, respectively. We demonstrate that FlexDeMo is faster than, and comparable in terms of validation loss, to both DeMo and extant hybrid sharding strategies that combine intra-node FSDP with full inter-node gradient synchronization. In essence, FlexDeMo enables training of larger DNNs more efficiently across multiple nodes of a cluster or even across geographically-dispersed clusters. We demonstrate that the Random replication-scheme, given the same bandwidth usage significantly outperforms DeMo and other replications schemes in translation, whereas we demonstrate DeMo to be superior in image-classification and causal language modeling. Last, we show that sharing the signed gradients is clearly beneficial.
@@
  • Comparable model performance compared to extant hybrid sharded training strategies using the AdamW optimizer.
@@
  • An analysis of the critical hyper-parameters TopK, chunk-size and sign providing guidance on its effects on efficiency and model performance.
@@
Strategies to both scale and accelerate the training of deep neural networks have been active research areas (Shoeybi et al. 2019; Rajbhandari et al. 2020) for quite some years. The importance of these methods effectiveness is increasing as we train more and more large language models both in and across growing HPC capacities. We will first clarify the terminology and then provide a brief overview of the most important advances in distributed training.
@@
The zero redundancy optimizer (ZeRO) (Rajbhandari et al. 2020), tackling data and model parallelism, has introduced a variety of strategies to train large models, including partitioning optimizer states, gradients, and parameters – all of which are being pulled dynamically to a single accelerator on demand. ZeRO has been integrated in the DeepSpeed library (Rasley et al. 2020). The DeepSpeed team has since iterated on ZeRO with releases of Zero-2 (DeepSpeed Team, Majumder, and Wang 2020) and Zero-3 (DeepSpeed Team 2021). While ZeRO aims to achieve memory-efficiency to enable training of very large models, it also comes with a severe communication overhead. Both DeMo (Peng, Quesnelle, and Kingma 2024) and our proposed FlexDeMo aim to alleviate this communication overhead.
FSDP (Zhao et al. 2023) takes inspiration from ZeRO, but makes a few modifications to the exchange of gradients: ZeRO uses a reduce-scatter operation to distribute the gradients and an all-gather operation for the updated parameters. FSDP instead uses an all-gather operation to recover unsharded parameters and a reduce-scatter operation for the gradients. FSDP also introduces a hybrid sharding strategy, enabling the user to define the set of sharding accelerators and corresponding replicates to allow for a trade-off between memory and throughput via sharding groups and gradient accumulation with or without communication.
Another common strategy is to optimize locally for multiple steps before synchronization (Stich 2019). DiLoCo (Douillard et al. 2023) employs federated averaging to enable training in poorly connected nodes through parallel local optimization and periodic global averaging.
Notably, there are several other advances in accelerating large scale neural network training, such as compressing the gradient by only considering its sign (Bernstein et al. 2018; Jiang et al. 2024) or through low-rank projection of the gradients (Zhao et al. 2024). Although tackling similar challenges, we consider these directions orthogonal to our work, as we are interested in improving the communication overhead regarding optimizer states.
@@ Methods
We first present our hybrid-sharded decoupled momentum optimizer, FlexDeMo, which allows divergent optimizer states across accelerator nodes. Sharding to the accelerators of a node and replicating between the nodes is a reasonable strategy. We introduce the Random and Striding replication schemes, compare with the priorly proposed DiLoCo, and challenge DeMo’s effectiveness both regarding convergence and computation. FlexDeMo can be used to shard and replicate between any sets of accelerators, though, allowing both replicating model shards to multiple subsets of accelerators on one node and distributing different shards across multiple nodes. Thereby, we are relaxing the model memory-constraint by sharding the model and optimizer states across multiple accelerators, typically within one node, it becomes possible to train significantly larger models. DeMo replication is guided by the fast moving components of the momentum, computed via the discrete cosine transform (DCT-II), $n$-strides indices for Striding and, naturally, randomly selected for Random.
One of the main practical limitations of DeMo is that it does not support FSDP or other sharding strategies at all and relies on an all_gather operation across multiple nodes, which is inherently slow, as it requires every training process to exchange and receive gradients from every other instance (see Appendix A). This raises the new research question regarding whether DeMo would be suitable in an sharded setting such as FSDP, as well as investigating if and how its hyperparameters need to be adjusted – in particular, how many fast-moving components to extract and further, if such components is most optimal to synchronize. Here, we set out to overcome these technical hurdles and investigate and validate the efficiency characteristics of such distributed training mechanism.
@@
During training, the optimizer states are communicated between the accelerators within the node. Then fast moving components are extracted following the method from DeMo (Peng, Quesnelle, and Kingma 2024). These components are compressed and exchanged between groups of accelerators. This method allows for training larger models, keeping the expensive communication within the groups of accelerators, and minimizing the expensive communication between groups.
@@
The flexible decoupled momentum training strategy we propose can be powered by multiple replication-schemes; The DeMo based Discrete Cosine Transform (DCT) of the momenta, Random, Striding and DiLoCo. The replication of the optimizer states is done between the groups of accelerators instead of between all accelerators and by only communicating the selected components (gradients). This is shown in the Appendix A, where accelerator 0 of node 0 is replicating momentum components to accelerator 0 of node 1, and so on. With this method, there is no replication of data between accelerator 0 in one node and accelerator 1 in another, effectively drastically limiting the amount of data shared between nodes.
@@
While we only sharing the selected gradients (e.g.: fast moving momentum components) between the nodes, and even only between accelerators with corresponding shards, we operate on the full gradients within each group, where the transfer speeds are usually significantly higher. This takes advantage of the typically high bandwidth within nodes, while acknowledging the slower bandwidth between nodes.
@@
To achieve this, we reimplement the stochastic gradient descent (SGD) with momentum optimizer from (Peng, Quesnelle, and Kingma 2024) (denoted DeMo-SGD), introducing a series of changes to support decoupled optimization in FSDP employing intra-node hybrid-sharding. We assume that the model is wrapped as an FSDP object. Both the forward and backward passes must be wrapped in a no_sync context manager, disabling automatic synchronization of gradients $\theta_{t}$ affecting the default all_reduce-functionality, consequently decoupling the momentum $m$ across accelerator-nodes. PyTorch’s Autograd produces the gradient for the whole unsharded grad-parameter accessible in p.grad.
@@
Shown in Appendix A is a schematic representation of the communication. We employ the reduce-scatter operation, averaging and then sharding the computed gradients back to their respective ranks (hence sharded parameter-size) in the sharding-parallel-group. This allow us to work only on the shards in the subsequent operations. We denote the two communication-groups as $S$ and $R$ referring to the sharding-group and replication-group, respectively. Gradients are scattered intra-node locally within $S$ and the fast components, $q$, are communicated inter-node in $R$. We describe the operation of FlexDeMo in Algorithm 1.
@@
The figure in Appendix A shows the communication pattern for standard DeMo within a single node, as well as the communication pattern for DeMo when extended to two nodes, with inter-node communication. DeMo does not employ any sharding, and is thus using an all_gather operation across the nodes, in the case of multiple nodes. This high amount of inter-node communication, is clearly evident in the figure.
@@
We are introducing the Random and Striding replication schemes and compare with the already proposed DiLoCo. Random denotes Random selection of indices. Striding denotes the selection of the n’th index and DiLoCo denotes synchronization each n’th optimization step. These replicators do not operate in chunks as the DeMo replicator. All replication schemes can achieve the same relative data compression. However, Striding and Random schemes do not need to share the selected indices with the gradient, as their operations can reproduced across training-instances with a fixed seed, enabling us to share double the amount of data, on the same bandwidth.
@@
AdamW is defacto the default optimizer for most objectives, prompting us to believe is also would demonstrate strong performance with these replication schemes. We implement AdamW, with reducing communication in mind, hence we do not share the first and seconds momenta, which would require 2-3 times more communication. The exponential moving average (EMA) and the moving average of the squared gradients are therefore not synchronized, and neither are the max of all exponential moving averages of squared gradients, if amsmax is employed.
@@
We present a series of experiments on FlexDeMo covering encoder-decoder language models, vision transformers, and decoder-only language models including scaling experiments, demonstraing the capabilties and training-behaviour across domains. We compare our Decoupled AdamW to DeMo-SGD as the underlying optimizer, challenging the DeMo replication-scheme with our introduced Striding and Random schemes, and further, compare with DiLoCo. Last, we demonstrate the performance and efficiency comparing performance to standard Hybrid-FSDP with a standard AdamW optimizer (Kingma and Ba 2015; Loshchilov and Hutter 2019) demonstrating the effectiveness of FlexDeMo.
Replication Schemes
Setup
First, we study hyperparameters for DeMo replication, including sign, topKK, transfer-dtype-type. Experiments can be found in the Appendix B. Here, we focus on small-scale experiments to study method design choices and hyperparameters. Specifically, we investigate the characteristics of the decoupled optimizations scheme in encoder-decoder architecture by training with a sequence-to-sequence objective. We train the T5-Large model on its native translation task on the English-to-French subset of the Opus Books Dataset (Tiedemann 2012).
The experiments are conducted on 2 nodes with 2 accelerators, employing Hybrid-FSDP. We run for 20 epochs, with a learning rate of $1e^{-3}$ and no weight decay. The optimizer and the replication scheme is depending on the specific experiment. These experiments are carried out on a HPC with a dragonfly network architecture with 200 Gbps interconnect. Nodes are NUMA-nodes and comprises of two GPU modules each with 4 AMD MI250x GPUs (128GB per module). The two modules are connected with infinity fabric connections (50+50 Gb/s) and the modules communicate over the network. We base our foundation experiments on the Opus Books dataset subset of English to French, standardizing our experimental setup for reproducibility.
Results: SGD vs. Decoupled AdamW
AdamW (Kingma and Ba 2015) is de-facto default optimizer, having demonstrated superior performance across many domains and tasks. However DeMo (Peng, Quesnelle, and Kingma 2024) is employing SGD (which we denote DeMo-SGD, we differentiate as it accumulates momenta.) as underlying optimizer. To investigate this design-choice and broaden the investigation, we set out to compare DeMo-SGD and our Decoupled-AdamW with employing DeMo, DiLoCo (Douillard et al. 2023), Random, and Striding replication schemes. We report validation loss in Figure 1. The experiments are conducted such that the bandwidth usage is identical in each setting (we denote this as the compression-rate), as it would be defined by hardware in practice. DeMo-SGD with Random replication demonstrates best performance. DeMo-SGD yields clearly better at generalization to the validation-set, with Random replication being best. Decoupled-AdamW and DeMo-SGD with DiLoCo replication following as seconds and third best, respectively. DeMo-SGD is improving overall as fast as the DiLoCo variants, but overall worse. Decoupled-AdamW with DeMo and Random replication schemes yields worst performance.
@@
We train a T5-base model on the French-to-English subset of Opus Books dataset (Tiedemann 2012) for 20 epochs, in a Hybrid-FSDP setup with 2 nodes with 8 accelerators each. We employ the SGD optimizer with Random, DeMo, DiLoCo and Striding replication. DeMo employs a chunk-size of 32. We hold out 20% as validation data.

### v4

@@ Abstract
Training large neural network models requires extensive computational resources, often distributed across several nodes and accelerators. Recent findings suggest that it may be sufficient to only exchange the fast-moving components of the gradients, while accumulating momentum locally (Decoupled Momentum, or DeMo). However, DeMo assumes that models fit on a single accelerator. We relax this assumption and introduce FlexDeMo, whereby nodes fully shard model parameters locally across different accelerators, while inter-node communication is reduced by synchronizing only fast-moving components instead of the full gradients – resulting in a hybrid sharded data parallel training strategy. We further introduce a framework, called DeToNATION, that generalizes DeMo, FlexDeMo, and other popular distributed training schemes such as DiLoCo – introducing new variations of replication schemes and challenging choices made in DeMo. Our results across language and vision domains show that FlexDeMo attains similar validation loss to hybrid sharded data parallel training employing AdamW and full gradient synchronization, while being substantially faster. FlexDeMo is thus a promising distributed training scheme for the largest machine learning models.
Codegithub.com/schneiderkamplab/DeToNATION/
@@ Introduction
Training large deep neural networks (DNNs) induces large amounts of network traffic in the form of gradients that are transmitted between accelerators, typically requiring expensive localized high-throughput network setups on high-performance computing clusters. The network throughput increasingly becomes a bottleneck, as the number of accelerator nodes participating in the training increases and the general network congestion increases. Recent work shows that synchronizing the full optimizer state is not always necessary for state-of-the-art results through decoupling momentum updates and allowing controlled divergence within each rank by carefully controlled inter-accelerator communication termed Decoupled Momentum optimization (DeMo) (peng2024demodecoupledmomentumoptimization).
DeMo presents a viable strategy for distributed training with reduced gradient communication and, thus, enables distributed data parallel (DDP) training of DNNs with relatively low network bandwidth requirements. However, this strategy comes with several caveats. First, relying on DDP implies the constraint that the DNN model and optimizer states must fit within the memory capacity of each accelerator. This severely limits the applicability for training large models such as state-of-the-art large language models, which commonly do not fit within the memory of a single accelerator. Second, DeMo inherently relies on a distributed gathering operation whose bandwidth requirements scale linearly with the number of accelerators. Third, DeMo leaves many open questions regarding the replication scheme and hyperparameters, such as chunk size, TopK and employing the sign-function.
@@
Instead of synchronizing the full gradients between the nodes as in extant hybrid sharding strategies, we compress, gather, and decompress selected relevant fast-moving momentum components following the approach introduced by DeMo for DDP training (peng2024demodecoupledmomentumoptimization). This relaxes constraints regarding the accelerator memory requirements while simultaneously reducing inter-node bandwidth requirements. We further investigate hyperparameters of DeMo and challenge the fast-moving momenta-based replication scheme by introducing Striding, Random and the previously proposed DiLoCo schemes. This addresses the three caveats of DeMo mentioned above. First, FlexDeMo allows for decoupled momentum training of models that do not fit into the memory of a single accelerator but fit into the combined memory of the accelerators of one node. Second, FlexDeMo replicates between nodes rather than accelerators, effectively reducing the bandwidth requirements of the distributed gathering operation, which now scales linearly with the number of nodes rather than the number of accelerators. Third, we justify our choice of hyperparameters by conducting experiments.
We validate our results on the T5, ViT, and OLMo models in machine translation, image classification and causal language modeling, respectively. We show that FlexDeMo is faster than both DeMo and extant hybrid sharding strategies while achieving comparable validation loss. In essence, FlexDeMo enables training of larger DNNs more efficiently across multiple nodes of a cluster or even across geographically-dispersed clusters. We show that the Random replication scheme, given the same bandwidth usage significantly outperforms DeMo and other replications schemes in translation, whereas we show DeMo to be superior in image-classification and causal language modeling. Finally, we show that sharing the signed gradients is clearly beneficial.
@@
  • Comparable model performance compared to hybrid sharded training strategies using the AdamW optimizer.
@@
  • An analysis of the critical hyper parameters TopK, chunk size and sign providing guidance on their effects on efficiency and model performance.
@@
Strategies to both scale and accelerate the training of deep neural networks have been active research areas (shoeybi2019megatron; rajbhandari2020zero) for several years. The importance of these methods’ effectiveness is increasing as we train more and more large language models both in and across growing HPC capacities. We will first clarify the terminology and then provide a brief overview of the most important advances in distributed training.
@@
The zero redundancy optimizer (ZeRO) (rajbhandari2020zero), tackling data and model parallelism, has introduced a variety of strategies to train large models, including partitioning optimizer states, gradients, and parameters – all of which are pulled dynamically to a single accelerator on demand. ZeRO has been integrated in the DeepSpeed library (rasley2020deepspeed). The DeepSpeed team has subsequently extended ZeRO with releases of Zero-2 (zero2) and Zero-3 (zero3). While ZeRO aims to achieve memory efficiency to enable training of very large models, it also comes with a substantial communication overhead. Both DeMo (peng2024demodecoupledmomentumoptimization) and our proposed FlexDeMo alleviate this communication overhead.
FSDP (zhao2023fsdp) takes inspiration from ZeRO but modifies how gradients are exchanged: ZeRO uses a reduce-scatter operation to distribute the gradients and an all-gather operation for the updated parameters. FSDP instead uses an all-gather operation to recover unsharded parameters and a reduce-scatter operation for the gradients. FSDP also introduces a hybrid sharding strategy, enabling users to define sharding groups to trade-off memory and throughput.
Another common strategy is to optimize locally for multiple steps before synchronization (DBLP:conf/iclr/Stich19). DiLoCo (douillard2023diloco) employs federated averaging to enable training in poorly connected nodes through parallel local optimization and periodic global averaging.
Notably, there are several other advances in accelerating large-scale neural network training, such as gradient compression by only considering its sign (bernstein2018signsgd; DBLP:conf/nips/JiangYYZ24) or through low-rank projection of the gradients (DBLP:conf/icml/Zhao0CWAT24). Although these approaches tackle similar challenges, we consider these directions orthogonal to our work, as we are interested in reducing the communication overhead regarding optimizer states.
@@ Methods
We first present our hybrid-sharded decoupled momentum optimizer, FlexDeMo, which allows divergent optimizer states across accelerator nodes. Sharding to the accelerators of a node and replicating between the nodes is a reasonable strategy. We introduce the Random and Striding replication schemes, compare with the previously proposed DiLoCo, and challenge DeMo’s effectiveness both regarding convergence and computation. FlexDeMo can be used to shard and replicate between any sets of accelerators, though, allowing both replicating model shards to multiple subsets of accelerators on one node and distributing different shards across multiple nodes. Thereby, we are relaxing the model memory constraint by sharding the model and optimizer states across multiple accelerators, typically within one node, it becomes possible to train significantly larger models. DeMo replication is guided by the fast moving components of the momentum, computed via the discrete cosine transform (DCT-II), $n$-stride indices for Striding and, naturally, randomly selected for Random.
One of the main practical limitations of DeMo is that it does not support FSDP or other sharding strategies at all and relies on an all_gather operation across multiple nodes, which is inherently slow, as it requires every training process to exchange and receive gradients from every other instance (see Appendix A). This raises the new research question regarding whether DeMo would be suitable in a sharded setting such as FSDP, as well as investigating if and how its hyperparameters need to be adjusted – in particular, how many fast-moving components to extract and further, if it is optimal to synchronize such components. Here, we overcome these technical hurdles and investigate and study efficiency characteristics of such distributed training mechanism.
@@
During training, the optimizer states are communicated between the accelerators within the node. Then fast moving components are extracted following the method from DeMo (peng2024demodecoupledmomentumoptimization). These components are compressed and exchanged between groups of accelerators. This method allows for training larger models, keeping the expensive communication within the groups of accelerators, and minimizing the expensive communication between groups.
@@
The flexible decoupled momentum training strategy we propose can be powered by multiple replication schemes; The DeMo based Discrete Cosine Transform (DCT) of the momenta, Random, Striding and DiLoCo. The replication of the optimizer states is done between the groups of accelerators instead of between all accelerators and by only communicating the selected components (gradients). That is, accelerator 0 of node 0 replicates momentum components to accelerator 0 of node 1, and so on (detail in Appendix A). With this method, there is no replication of data between accelerator 0 in one node and accelerator 1 in another, drastically limiting the amount of data shared between nodes.
@@
While we only share the selected gradients (e.g.: fast moving momentum components) between the nodes, and even only between accelerators with corresponding shards, we operate on all gradients within each group, where the transfer speeds are usually significantly higher. This takes advantage of the typically high bandwidth within nodes, while acknowledging the lower bandwidth between nodes.
@@
To achieve this, we reimplement the stochastic gradient descent (SGD) with momentum optimizer from (peng2024demodecoupledmomentumoptimization) (denoted DeMo-SGD), introducing a series of changes to support decoupled optimization in FSDP employing intra-node hybrid sharding. We assume that the model is wrapped as an FSDP object. Both the forward and backward passes must be wrapped in a no_sync context manager, disabling automatic synchronization of gradients $\theta_{t}$, affecting the default all_reduce-functionality, consequently decoupling the momentum $m$ across accelerator-nodes. PyTorch’s Autograd produces the gradient for the whole unsharded grad-parameter accessible in p.grad.
@@
We employ the reduce-scatter operation, averaging and then sharding the computed gradients back to their respective ranks (hence sharded parameter-size) in the sharding-parallel-group. This allows us to work only on the shards in the subsequent operations. We denote the two communication-groups as $S$ and $R$ referring to the sharding-group and replication group, respectively. Gradients are scattered intra-node locally within $S$ and the fast components, $q$, are communicated inter-node in $R$. We describe the operation of FlexDeMo in Algorithm 1 and provide a diagram of the communication in Appendix A.
@@
A major difference to pure DeMo, is that it does not employ any sharding, and is thus using an all_gather operation across the nodes, in the case of multiple nodes. This high amount of inter-node communication, is clearly evident in the figure in Appendix A.
@@
We are introducing the Random and Striding replication schemes and compare with the already proposed DiLoCo. Random replication denotes a random selection of $n$ indices. Striding denotes the selection of every $n$’th index, and DiLoCo denotes synchronization every n’th optimization step. These replicators do not operate on chunks of the gradients as the DeMo replicator. All replication schemes can achieve the same relative data compression. However, on top of the limited amount of gradient values shared, Striding and Random schemes do not need to share the selected indices of the gradients, as they can be reproduced across training-instances with a fixed seed, effectively halving the data transfer, for the same amount of shared gradients.
@@
AdamW is defacto the default optimizer for most objectives, prompting us to believe that it would demonstrate strong performance with these replication schemes. We implement AdamW without synchronizing the first and second momenta, which would require communicating 2-3 times more data. The exponential moving average (EMA) and the moving average of the squared gradients are therefore not synchronized, and neither are the max of all exponential moving averages of squared gradients, if amsmax is employed.
@@
We present a series of experiments on FlexDeMo covering encoder-decoder language models, vision transformers, and decoder-only language models including scaling experiments, demonstrating the capabilties and training behavior across domains. We compare our Decoupled AdamW to DeMo-SGD as the underlying optimizer, challenging the DeMo replication scheme with our introduced Striding and Random schemes, and further, compare with DiLoCo. Last, we demonstrate the performance and efficiency comparing performance to standard Hybrid-FSDP with a standard AdamW optimizer (kingma2014adam; adamw) demonstrating the effectiveness of FlexDeMo.
Overall Setup
First, we study hyperparameters for DeMo replication, including sign, TopK, transfer-dtype. Experiments can be found in Appendix B. Summarized, all replicators show that only sharing the sign of the gradients is beneficial for the loss convergence. It is also clear that fp32 converges faster than fp16 for the dtype, again for all replication schemes. As for TopK and chunk size, the conclusion is less clear. For some cases, larger chunk sizes converge faster, while for other the optimum is lower chunk sizes. For all experiments conducted, we choose $\mathrm{chunksize}=32$, dtype=fp32, and syncing only the sign of the gradients. TopK is varied through the experiments, to control the compression ratio, given by $\frac{\mathrm{topK}}{\mathrm{chunksize}}$. The compression rate is the percentage of gradients shared between nodes.
In the upcoming experiments we will demonstrate the performance across sequence-to-sequence, image-classification and causal language modeling, using the T5-base, ViT-B and OLMo2 models. The experiments are carried out on a HPC with a dragonfly network architecture with 200 Gbps interconnect. Nodes are NUMA-nodes and comprised of two GPU modules each with 4 AMD MI250x GPUs (128GB per module). The two modules are connected with infinity fabric connections (50+50 Gb/s) through which they communicate.
@@
We train a T5-base model on the French-to-English subset of Opus Books dataset (tiedemann-2012-parallel). We employ the SGD optimizer at a learning rate of $10^{-3}$, with Random, DeMo, DiLoCo and Striding replication. We hold out 20% as validation data.